The Boeing Company (BA)

NYSE - NYSE Delayed Price. Currency in USD

362.17-10.53 (-2.83%)

At close: 4:00PM EDT

361.65 -0.52 (-0.14%)
After hours: 7:57PM EDT

Buy

Sel

Previous Close

372.70

Open

368.79

Bid

361.50 x 1200

Ask

362.34 x 2900

Day's Range

361.52 - 369.86

52 Week Range

292.47 - 446.01

Market Cap

?????

Beta (3Y Monthly)

1.36

PE Ratio (TTM)

20.29

EPS (TTM)

?????

Forward Dividend & Yield

8.22 (???%)

(a) BA had $10.4 Billion of net income in the recent year. What is the BA’s market cap value? Answer the question using the information above.

(c) What is the BA’s dividend yield for the next quarter? Answer the question using the information above.

(d) What is the expected stock price next year if the current PE ratio (TTM) is used as a benchmark and EPS(TTM) is expected to grow 3% next year?
Answer #1

1.
=Net Income*PE ratio
=10.4*20.29=211.02 billion

2.
=Dividend/Price=8.22/362.17=2.270%

3.
=Price*(1+growth rate)
=362.17*(1+3%)=$373.035100